# L3Cube-IndicNews: News-based Short Text and Long Document Classification Datasets in Indic Languages

### v1

@@ Abstract
In this work, we introduce L3Cube-IndicNews, a multilingual text classification corpus aimed at curating a high-quality dataset for Indian regional languages, with a specific focus on news headlines and articles. We have centered our work on 10 prominent Indic languages, including Hindi, Bengali, Marathi, Telugu, Tamil, Gujarati, Kannada, Odia, Malayalam, and Punjabi. Each of these news datasets comprises 10 or more classes of news articles. L3Cube-IndicNews offers 3 distinct datasets tailored to handle different document lengths that are classified as: Short Headlines Classification (SHC) dataset containing the news headline and news category, Long Document Classification (LDC) dataset containing the whole news article and the news category, and Long Paragraph Classification (LPC) containing sub-articles of the news and the news category. We maintain consistent labeling across all 3 datasets for in-depth length-based analysis. We evaluate each of these Indic language datasets using 4 different models including monolingual BERT, multilingual Indic Sentence BERT (IndicSBERT), and IndicBERT. This research contributes significantly to expanding the pool of available text classification datasets and also makes it possible to develop topic classification models for Indian regional languages. This also serves as an excellent resource for cross-lingual analysis owing to the high overlap of labels among languages. The datasets and models are shared publicly at https://github.com/l3cube-pune/indic-nlp.
@@
[…]
@@ 1 Introduction
India boasts a rich linguistic diversity, with over 700 languages spoken, out of which 22 are officially recognized. Some of the primarily spoken languages include Hindi, Bengali, Marathi, Telugu, Tamil, Gujarati, Kannada, Odia, Malayalam, and Punjabi. Despite their widespread use, there’s a notable scarcity of comprehensive Indic language datasets, primarily due to their low-resource status and linguistic complexity Patil and Patil (2017).
@@
  • The datasets are bench-marked using state-of-the-art pre-trained BERT models: L3Cube monolingual BERT, L3Cube monolingual SBERT, L3Cube IndicSBERT (multilingual) and IndicBERT.
@@ 2 Related work
IndicNLP News Article Classification333 https://github.com/AI4Bharat/indicnlp_corpus#indicnlp-news-article-classification-dataset dataset is part of the AI4Bharat-IndicNLP Dataset Kunchukuttan et al. (2020) that consists of news articles in 10 Indian languages categorized into classes like sports, entertainment, business, politics, and lifestyle. While this dataset contains a substantial number of records, it falls short in terms of the variety of categories available for news articles in each language, thereby limiting its diversity.
Multi Indic Languages News Dataset444 https://www.kaggle.com/datasets/shaz13/multi-indic-languages-news-datasethttps://www.kaggle.com/datasets/shaz13/multi-indic-languages-news-dataset is a dataset publicly available on Kaggle. It is a multi-language news dataset from Times Internet for various Indian languages. This data contains columns named title, link, description, long_description, id. Despite the extensive size and diversity of this dataset, encompassing a rich collection of records across numerous Indian languages, it lacks language-wise and news category-wise segregation, which hampers clarity and ease of use.
@@
iNLTK666 https://github.com/goru001/inltk is an openly accessible dataset primarily comprising data for 13 Indic languages, sourced from Wikipedia articles. It encompasses over 12,000 cleaned rows for each of these languages Arora (2020). This dataset comprises publicly accessible data for languages like Hindi, Bengali, Punjabi, Kannada, and Oriya. For languages such as Gujarati, Malayalam, Marathi, and Tamil, they have created their dataset by extracting information from Wikipedia articles. While the dataset boasts a substantial volume of records for each language, it falls short in terms of categorizing the data into specific news categories. Furthermore, in some languages, it includes not only news-related articles but also other types of content, leading to inefficiency and inconsistency in the dataset’s content and structure.
@@
Hindi BBC Articles
India, International, Entertainment, Sports, Others
172K
BBC Hindi News Articles
India, Pakistan, News, International, Sntertainment, port, Science, China, Learningenglish, Social, Southasia, Business, Institutional, Multimedia
4335
Bengali AI4Bharat-IndicNLP Dataset
Entertainment, Sports
14K
Soham Articles
Kolkata, State, National, International, Sports, Entertainment
72K
Marathi AI4Bharat-IndicNLP Dataset
Entertainment, Lifestyle, Sports
4.5K
iNLTK Headlines
State, Entertainment, Sports
85K
Telugu AI4Bharat-IndicNLP Dataset
Entertainment, Business, Sports
24K
Tamil AI4Bharat-IndicNLP Dataset
Entertainment, Politics, Sports
11.7K
iNLTK Headlines
Tamil-cinema, Business, Spirituality
127K
Gujarati AI4Bharat-IndicNLP Dataset
Business, Entertainment, Sports
2K
iNLTK Headlines
Entertainment, Business, Tech
31K
Kannada AI4Bharat-IndicNLP Dataset
Entertainment, Lifestyle, Sports
30K
iNLTK-IndicNLP News Category
Entertainment, Sports, Tech
6.3K
Odia AI4Bharat-IndicNLP Dataset
Business, Crime, Entertainment, Sports
30K
iNLTK-IndicNLP News Category
Sports, Business, Entertainment
19K
Malayalam AI4Bharat-IndicNLP Dataset
Business, Entertainment, Sports, Technology
6K
iNLTK Headlines
Entertainment, Sports, Business
12K
Punjabi AI4Bharat-IndicNLP Dataset
Business, Crime, Entertainment, sports
3.1K
iNLTK -IndicNLP News Category
Politics, Non-politics
800
@@ 3.1 Data collection
For the dataset curation process, we identified several websites to collect a substantial number of articles for each news category. The Hindi language dataset was scraped from Jansatta777 https://www.jansatta.com/. The Marathi language dataset was scraped from Lokmat888 https://www.lokmat.com/. The Bengali language dataset was scraped from Aajkal999 https://www.aajkaal.in/, Ganashakti101010 https://ganashakti.com/, BBC111111 https://www.bbc.com/bengali, Anandabazar121212 https://www.anandabazar.com/, abnews24131313 https://www.sangbadpratidin.in/, and Sangbadpratidin141414 https://www.abnews24.com/. The Telugu language dataset was scraped from ABP Telugu151515 https://telugu.abplive.com/. The Tamil Language dataset was scraped from Hindu Tamil161616 https://www.hindutamil.in/. The Gujarati language dataset was scraped from ABP Gujarati171717 https://gujarati.abplive.com/. The Kannada language dataset was scraped from Kannada Prabha181818 https://www.kannadaprabha.com/ and PublicTV191919 https://publictv.in/. The Odia language dataset was scraped from Odisha Bhaskar202020 https://odishabhaskar.com/ and Dharitri212121 https://www.dharitri.com/https://www.dharitri.com/ . The Malayalam language dataset was scraped from Madhyamam222222 https://www.madhyamam.com/. The Punjabi language dataset was scraped from Khabarwaale232323 https://www.khabarwaale.com/.
@@ 3.3 Dataset Statistics
Hindi (11)
Auto, Business, Crime, Education, Entertainment, Health, International, Nation, Politics, Sports, Technology
Bengali (10)
Sports, National, International, Kolkata, State, Politics, Entertainment, Technology, Editorial, Lifestyle
Marathi (12)
Auto, Bhakti, Crime, Education, Fashion, Health, International, Manoranjan, Politics, Sports, Tech, Travel
Telugu (10)
Business, Crime, Education, Entertainment, Jobs, Lifestyle, Politics, Sports, Technology, World
Tamil (10)
Auto, Business, Crime, Education, Entertainment, Health, India, Lifestyle, Politics, World
Gujarati (10)
Astro, Auto, Business, Crime, Entertainment, International, Nation, Sports, State, Technology
Kannada (10)
Business, Crime, Cuisine, Entertainment, International, Nation, Politics, Sports, State, Technology
Odia (10)
Lifestyle, Entertainment, News, Crime, Business, Health, Politics, Career, Sports, Editorial
Malayalam (10)
Crime, Entertainment, Gulf, International, Kerala, Lifestyle, National, Opinion, Sports, Technology
Punjabi (10)
World, Sports, Religious, Education, Transfer & Appointments, Literature, Health, National, Crime, Lifestyle
@@ 4 Models
4.1 L3Cube Monolingual BERT242424 https://arxiv.org/abs/2211.11418 for Indic languages
We use the monolingual BERT models for the 10 Indic languages, released by L3cube-Pune252525 https://github.com/l3cube-pune as the base models. These models are termed as HindBERT262626 https://huggingface.co/l3cube-pune/hindi-bert-v2, BengaliBERT272727 https://huggingface.co/l3cube-pune/bengali-bert, MahaBERT282828 https://huggingface.co/l3cube-pune/marathi-bert, TeluguBERT292929 https://huggingface.co/l3cube-pune/telugu-bert, TamilBERT303030 https://huggingface.co/l3cube-pune/tamil-bert, GujaratiBERT313131 https://huggingface.co/l3cube-pune/gujarati-bert, KannadaBERT323232 https://huggingface.co/l3cube-pune/kannada-bert, OdiaBERT333333 https://huggingface.co/l3cube-pune/odia-bert, MalayalamBERT343434 https://huggingface.co/l3cube-pune/malayalam-bert, PunjabiBERT353535 https://huggingface.co/l3cube-pune/punjabi-bert. These models are fine-tuned on the existing multilingual models like MuRIL Khanuja et al. (2021), xlmRoBERTa Conneau et al. (2019), and IndicBERT on the monolingual corpus.
@@
We also evaluate L3Cube monolingual Indic SBERT models that are HindSBERT373737 https://huggingface.co/l3cube-pune/hindi-sentence-bert-nliJoshi et al. (2023b), BengaliSBERT383838 https://huggingface.co/l3cube-pune/bengali-sentence-bert-nli, MahaSBERT [^39] https://huggingface.co/l3cube-pune/marathi-sentence-bert-nli, TeluguSBERT404040 https://huggingface.co/l3cube-pune/telugu-sentence-bert-nli, TamilSBERT414141 https://huggingface.co/l3cube-pune/tamil-sentence-bert-nli, GujaratiSBERT424242 https://huggingface.co/l3cube-pune/gujarati-sentence-bert-nli, KannadaSBERT434343 https://huggingface.co/l3cube-pune/kannada-sentence-bert-nli, OdiaSBERT444444 https://huggingface.co/l3cube-pune/odia-sentence-bert-nli, MalayalamSBERT454545 https://huggingface.co/l3cube-pune/malayalam-sentence-bert-nli, and PunjabiSBERT464646 https://huggingface.co/l3cube-pune/punjabi-sentence-bert-nli.
@@ 5 Evaluation
HindiBERT
86.600 91.681 88.097
Hindi
HindiSBERT
86.518
91.534
87.810
IndicSBERT
85.870
91.604
86.922
IndicBERT
83.910
86.809
84.424
BengaliBERT
82.549
95.293 85.195
Bengali
BengaliSBERT
82.749
94.141
84.470
IndicSBERT
80.674
92.457
85.232
IndicBERT
80.007
90.757
82.457
MarathiBERT
91.163 94.706
86.731
Marathi
MarathiSBERT
91.017
94.349
87.439
IndicSBERT
90.510
93.987
87.103
IndicBERT
89.388
92.627
85.222
TeluguBERT
89.810
92.765 91.818
Telugu
TeluguSBERT
90.416
92.651
91.098
IndicSBERT
87.916
92.348
91.515
IndicBERT
88.371
92.943
91.811
TamilBERT
81.785
84.521
81.300
Tamil
TamilSBERT
81.720
86.122
79.573
IndicSBERT
81.209
84.227
79.703
IndicBERT
81.275
83.226
80.571
GujaratiBERT
89.898 95.278 90.640
Gujarati
GujaratiSBERT
89.808
95.060
90.091
IndicSBERT
88.613
95.277
89.573
IndicBERT
87.909
90.854
89.050
KannadaBERT
91.410
94.706
87.704
Kannada
KannadaSBERT
89.340
94.539
87.345
IndicSBERT
90.255
94.833
87.835
IndicBERT
87.965
91.857
87.358
OdiaBERT
83.399
92.940
84.557
Odia
OdiaSBERT
84.137 93.424
83.918
IndicSBERT
82.065
93.054
84.441
IndicBERT
82.661
86.962
82.322
MalayalamBERT
80.440 88.573 81.705
Malayalam
MalayalamSBERT
80.171
88.201
81.171
IndicSBERT
77.780
88.029
79.672
IndicBERT
75.114
83.098
76.457
PunjabiBERT
85.456
90.182
84.163
Punjabi
PunjabiSBERT
90.363
94.866
90.333
IndicSBERT
89.490
94.224
88.446
IndicBERT
91.725
94.781
90.358
@@
Hindi
HindiSBERT
86.209
88.451
88.279
Bengali
BengaliSBERT
82.615
93.126
89.289
Marathi
MarathiSBERT
91.017
94.349
87.439
Telugu
TeluguSBERT
88.674
92.310
91.174
Tamil
TamilSBERT
81.754
84.902
80.993
Gujarati
GujaratiSBERT
88.792
92.378
90.609
Kannada
KannadaSBERT
89.797
94.212
88.260
Odia
OdiaSBERT
84.137
93.424
83.918
Malayalam
MalayalamSBERT
79.114
87.743
80.428
Punjabi
PunjabiSBERT
94.744
93.614
93.782

### v2

@@ Abstract
In this work, we introduce L3Cube-IndicNews, a multilingual text classification corpus aimed at curating a high-quality dataset for Indian regional languages, with a specific focus on news headlines and articles. We have centered our work on 11 prominent Indic languages, including Hindi, Bengali, Marathi, Telugu, Tamil, Gujarati, Kannada, Odia, Malayalam, Punjabi and English. Each of these news datasets comprises 10 or more classes of news articles. L3Cube-IndicNews offers 3 distinct datasets tailored to handle different document lengths that are classified as: Short Headlines Classification (SHC) dataset containing the news headline and news category, Long Document Classification (LDC) dataset containing the whole news article and the news category, and Long Paragraph Classification (LPC) containing sub-articles of the news and the news category. We maintain consistent labeling across all 3 datasets for in-depth length-based analysis. We evaluate each of these Indic language datasets using 4 different models including monolingual BERT, multilingual Indic Sentence BERT (IndicSBERT), and IndicBERT. This research contributes significantly to expanding the pool of available text classification datasets and also makes it possible to develop topic classification models for Indian regional languages. This also serves as an excellent resource for cross-lingual analysis owing to the high overlap of labels among languages. The datasets and models are shared publicly at https://github.com/l3cube-pune/indic-nlp.
@@
Aishwarya Mirashi1,3, Srushti Sonavane1,3, Purva Lingayat1,3, Tejas Padhiyar1,3and Raviraj Joshi2,3 Pune Institute of Computer Technology, Pune1 Indian Institute of Technology Madras, Chennai2 L3Cube Labs, Pune3
@@ 1 Introduction
India boasts a rich linguistic diversity, with over 700 languages spoken, out of which 22 are officially recognized. Some of the primarily spoken languages include Hindi, Bengali, Marathi, Telugu, Tamil, Gujarati, Kannada, Odia, Malayalam, Punjabi and English. Despite their widespread use, there’s a notable scarcity of comprehensive Indic language datasets, primarily due to their low-resource status and linguistic complexity Patil and Patil (2017).
@@
  • The datasets are bench-marked using state-of-the-art pre-trained BERT models: L3Cube monolingual BERT, L3Cube monolingual SBERT, L3Cube IndicSBERT (multilingual) and IndicBERT. The models for individual languages are shared publicly on Hugging Face [^3] https://huggingface.co/l3cube-pune/marathi-topic-all-doc-v2(see Appendix).
@@ 2 Related work
IndicNLP News Article Classification444 https://github.com/AI4Bharat/indicnlp_corpus#indicnlp-news-article-classification-dataset dataset is part of the AI4Bharat-IndicNLP Dataset Kunchukuttan et al. (2020) that consists of news articles in 10 Indian languages categorized into classes like sports, entertainment, business, politics, and lifestyle. While this dataset contains a substantial number of records, it falls short in terms of the variety of categories available for news articles in each language, thereby limiting its diversity.
Multi Indic Languages News Dataset555 https://www.kaggle.com/datasets/shaz13/multi-indic-languages-news-datasethttps://www.kaggle.com/datasets/shaz13/multi-indic-languages-news-dataset is a dataset publicly available on Kaggle. It is a multi-language news dataset from Times Internet for various Indian languages. This data contains columns named title, link, description, long_description, id. Despite the extensive size and diversity of this dataset, encompassing a rich collection of records across numerous Indian languages, it lacks language-wise and news category-wise segregation, which hampers clarity and ease of use.
@@
iNLTK777 https://github.com/goru001/inltk is an openly accessible dataset primarily comprising data for 13 Indic languages, sourced from Wikipedia articles. It encompasses over 12,000 cleaned rows for each of these languages Arora (2020). This dataset comprises publicly accessible data for languages like Hindi, Bengali, Punjabi, Kannada, and Oriya. For languages such as Gujarati, Malayalam, Marathi, and Tamil, they have created their dataset by extracting information from Wikipedia articles. While the dataset boasts a substantial volume of records for each language, it falls short in terms of categorizing the data into specific news categories. Furthermore, in some languages, it includes not only news-related articles but also other types of content, leading to inefficiency and inconsistency in the dataset’s content and structure.
@@
Hindi BBC Articles India, International, Entertainment, Sports, Others 172K
BBC Hindi News Articles India, Pakistan, News, International, Sntertainment, port, Science, China, Learningenglish, Social, Southasia, Business, Institutional, Multimedia 4335
Bengali AI4Bharat-IndicNLP Dataset Entertainment, Sports 14K
Soham Articles Kolkata, State, National, International, Sports, Entertainment 72K
Marathi AI4Bharat-IndicNLP Dataset Entertainment, Lifestyle, Sports 4.5K
iNLTK Headlines State, Entertainment, Sports 85K
Telugu AI4Bharat-IndicNLP Dataset Entertainment, Business, Sports 24K
Tamil AI4Bharat-IndicNLP Dataset Entertainment, Politics, Sports 11.7K
iNLTK Headlines Tamil-cinema, Business, Spirituality 127K
Gujarati AI4Bharat-IndicNLP Dataset Business, Entertainment, Sports 2K
iNLTK Headlines Entertainment, Business, Tech 31K
Kannada AI4Bharat-IndicNLP Dataset Entertainment, Lifestyle, Sports 30K
iNLTK-IndicNLP News Category Entertainment, Sports, Tech 6.3K
Odia AI4Bharat-IndicNLP Dataset Business, Crime, Entertainment, Sports 30K
iNLTK-IndicNLP News Category Sports, Business, Entertainment 19K
Malayalam AI4Bharat-IndicNLP Dataset Business, Entertainment, Sports, Technology 6K
iNLTK Headlines Entertainment, Sports, Business 12K
Punjabi AI4Bharat-IndicNLP Dataset Business, Crime, Entertainment, sports 3.1K
iNLTK -IndicNLP News Category Politics, Non-politics 800
@@ 3.1 Data collection
For the dataset curation process, we identified several websites to collect a substantial number of articles for each news category. The Hindi language dataset was scraped from Jansatta888 https://www.jansatta.com/. The Marathi language dataset was scraped from Lokmat999 https://www.lokmat.com/. The Bengali language dataset was scraped from Aajkal101010 https://www.aajkaal.in/, Ganashakti111111 https://ganashakti.com/, BBC121212 https://www.bbc.com/bengali, Anandabazar131313 https://www.anandabazar.com/, abnews24141414 https://www.sangbadpratidin.in/, and Sangbadpratidin151515 https://www.abnews24.com/. The Telugu language dataset was scraped from ABP Telugu161616 https://telugu.abplive.com/. The Tamil Language dataset was scraped from Hindu Tamil171717 https://www.hindutamil.in/. The Gujarati language dataset was scraped from ABP Gujarati181818 https://gujarati.abplive.com/. The Kannada language dataset was scraped from Kannada Prabha191919 https://www.kannadaprabha.com/ and PublicTV202020 https://publictv.in/. The Odia language dataset was scraped from Odisha Bhaskar212121 https://odishabhaskar.com/ and Dharitri222222 https://www.dharitri.com/https://www.dharitri.com/ . The Malayalam language dataset was scraped from Madhyamam232323 https://www.madhyamam.com/. The Punjabi language dataset was scraped from Khabarwaale242424 https://www.khabarwaale.com/.
@@ 3.3 Dataset Statistics
Hindi (11) Auto, Business, Crime, Education, Entertainment, Health, International, Nation, Politics, Sports, Technology
Bengali (10) Sports, National, International, Kolkata, State, Politics, Entertainment, Technology, Editorial, Lifestyle
Marathi (12) Auto, Bhakti, Crime, Education, Fashion, Health, International, Manoranjan, Politics, Sports, Tech, Travel
Telugu (10) Business, Crime, Education, Entertainment, Jobs, Lifestyle, Politics, Sports, Technology, World
Tamil (10) Auto, Business, Crime, Education, Entertainment, Health, India, Lifestyle, Politics, World
Gujarati (10) Astro, Auto, Business, Crime, Entertainment, International, Nation, Sports, State, Technology
Kannada (10) Business, Crime, Cuisine, Entertainment, International, Nation, Politics, Sports, State, Technology
Odia (10) Lifestyle, Entertainment, News, Crime, Business, Health, Politics, Career, Sports, Editorial
Malayalam (10) Crime, Entertainment, Gulf, International, Kerala, Lifestyle, National, Opinion, Sports, Technology
Punjabi (10) World, Sports, Religious, Education, Transfer & Appointments, Literature, Health, National, Crime, Lifestyle
English (10) Health, Business, Elections, Education, Lifestyle, World, Sports, Entertainment, Science, Auto
@@ 4 Models
4.1 L3Cube Monolingual BERT252525 https://arxiv.org/abs/2211.11418 for Indic languages
We use the monolingual BERT models for the 10 Indic languages, released by L3cube-Pune262626 https://github.com/l3cube-pune as the base models. These models are termed as HindBERT272727 https://huggingface.co/l3cube-pune/hindi-bert-v2, BengaliBERT282828 https://huggingface.co/l3cube-pune/bengali-bert, MahaBERT292929 https://huggingface.co/l3cube-pune/marathi-bert, TeluguBERT303030 https://huggingface.co/l3cube-pune/telugu-bert, TamilBERT313131 https://huggingface.co/l3cube-pune/tamil-bert, GujaratiBERT323232 https://huggingface.co/l3cube-pune/gujarati-bert, KannadaBERT333333 https://huggingface.co/l3cube-pune/kannada-bert, OdiaBERT343434 https://huggingface.co/l3cube-pune/odia-bert, MalayalamBERT353535 https://huggingface.co/l3cube-pune/malayalam-bert, PunjabiBERT363636 https://huggingface.co/l3cube-pune/punjabi-bert. These models are fine-tuned on the existing multilingual models like MuRIL Khanuja et al. (2021), xlmRoBERTa Conneau et al. (2019), and IndicBERT on the monolingual corpus.
@@
We also evaluate L3Cube monolingual Indic SBERT models that are HindSBERT383838 https://huggingface.co/l3cube-pune/hindi-sentence-bert-nliJoshi et al. (2023b), BengaliSBERT393939 https://huggingface.co/l3cube-pune/bengali-sentence-bert-nli, MahaSBERT [^40] https://huggingface.co/l3cube-pune/marathi-sentence-bert-nli, TeluguSBERT414141 https://huggingface.co/l3cube-pune/telugu-sentence-bert-nli, TamilSBERT424242 https://huggingface.co/l3cube-pune/tamil-sentence-bert-nli, GujaratiSBERT434343 https://huggingface.co/l3cube-pune/gujarati-sentence-bert-nli, KannadaSBERT444444 https://huggingface.co/l3cube-pune/kannada-sentence-bert-nli, OdiaSBERT454545 https://huggingface.co/l3cube-pune/odia-sentence-bert-nli, MalayalamSBERT464646 https://huggingface.co/l3cube-pune/malayalam-sentence-bert-nli, and PunjabiSBERT474747 https://huggingface.co/l3cube-pune/punjabi-sentence-bert-nli.
@@ 5 Evaluation
HindiBERT 86.600 91.681 88.097
Hindi HindiSBERT 86.518 91.534 87.810
IndicSBERT 85.870 91.604 86.922
IndicBERT 83.910 86.809 84.424
BengaliBERT 82.549 95.293 85.195
Bengali BengaliSBERT 82.749 94.141 84.470
IndicSBERT 80.674 92.457 85.232
IndicBERT 80.007 90.757 82.457
MarathiBERT 91.163 94.706 86.731
Marathi MarathiSBERT 91.017 94.349 87.439
IndicSBERT 90.510 93.987 87.103
IndicBERT 89.388 92.627 85.222
TeluguBERT 89.810 92.765 91.818
Telugu TeluguSBERT 90.416 92.651 91.098
IndicSBERT 87.916 92.348 91.515
IndicBERT 88.371 92.943 91.811
TamilBERT 81.785 84.521 81.300
Tamil TamilSBERT 81.720 86.122 79.573
IndicSBERT 81.209 84.227 79.703
IndicBERT 81.275 83.226 80.571
GujaratiBERT 89.898 95.278 90.640
Gujarati GujaratiSBERT 89.808 95.060 90.091
IndicSBERT 88.613 95.277 89.573
IndicBERT 87.909 90.854 89.050
KannadaBERT 91.410 94.706 87.704
Kannada KannadaSBERT 89.340 94.539 87.345
IndicSBERT 90.255 94.833 87.835
IndicBERT 87.965 91.857 87.358
OdiaBERT 83.399 92.940 84.557
Odia OdiaSBERT 84.137 93.424 83.918
IndicSBERT 82.065 93.054 84.441
IndicBERT 82.661 86.962 82.322
MalayalamBERT 80.440 88.573 81.705
Malayalam MalayalamSBERT 80.171 88.201 81.171
IndicSBERT 77.780 88.029 79.672
IndicBERT 75.114 83.098 76.457
PunjabiBERT 85.456 90.182 84.163
Punjabi PunjabiSBERT 90.363 94.866 90.333
IndicSBERT 89.490 94.224 88.446
IndicBERT 91.725 94.781 90.358
Bert-base uncased 76.634 92.177 78.392
English IndicSBERT 75.237 91.070 77.142
IndicBERT 76.788 90.813 77.913
@@
Hindi HindiSBERT 86.209 88.451 88.279
Bengali BengaliSBERT 82.615 93.126 89.289
Marathi MarathiSBERT 91.017 94.349 87.439
Telugu TeluguSBERT 88.674 92.310 91.174
Tamil TamilSBERT 81.754 84.902 80.993
Gujarati GujaratiSBERT 88.792 92.378 90.609
Kannada KannadaSBERT 89.797 94.212 88.260
Odia OdiaSBERT 84.137 93.424 83.918
Malayalam MalayalamSBERT 79.114 87.743 80.428
Punjabi PunjabiSBERT 94.744 93.614 93.782
English Bert-base 74.28 94.87 75.37